OpenStudy (anonymous):

choose 2 fractions between -13/15 & -12/15

OpenStudy (anonymous):

-25/30 -51/60

OpenStudy (mertsj):

-13/15 = -52/60 -12/15 = -48/60 Fractions in between are -49/60, -50/60, -51/60 Take your pick

OpenStudy (anonymous):

how 'bout this: -51/60 ; -128/150

OpenStudy (anonymous):

r they between -13/15 & -12/15?

OpenStudy (lilg132):

dont bother helping him doesnt even click good answer